# Gravitational wave probes of particle dark matter: a review

### v3

@@ I Introduction
Our knowledge of physics has been incomplete for decades, if not for over a century Bertone and Hooper (2018). Dark energy and dark matter (DM) could comprise 69% and 26% of the universe, and yet we have not been able to directly detect either. The experimental evidence for the existence of some “invisible” or DM is overwhelming: stars orbiting the center of the Milky Way would have flown out of their orbits if DM did not hold them in place Freeman (1970); Rogstad and Shostak (1972); Whitehurst and Roberts (1972); Roberts and Rots (1973); anisotropies in the cosmic microwave background could result from DM potential wells in the early universe Peebles (1982); Burles and Tytler (1998); DM could seed the formation of large-scale structures visible today, which would have been impossible with ordinary matter alone Pagels and Primack (1982); Davis et al. (1982); Blumenthal et al. (1984); Davis et al. (1985); Navarro et al. (1996); and gravitational lensing of light by galaxies in between us and the light source could be explained by invisible mass inside the galaxy Paczynski (1986); Alcock et al. (1993); Aubourg et al. (1993); Alcock et al. (2000); Lasserre (2000); Tisserand et al. (2007).
@@
While the observational evidence indicates that DM may exist, it does not explain what it is made of. Could it be a single new particle, such as the axion Peccei and Quinn (1977a, b); Weinberg (1978) or dark photon Galison and Manohar (1984); Holdom (1986); Pierce et al. (2018); Filippi and De Napoli (2020), a collection of new particles, or macroscopic, in the form of PBHs formed within $\sim 1$ second of the Big Bang Green and Kavanagh (2021)? Or, particle DM may (or may not) exist alongside PBHs, which may comprise a portion or the totality of DM Hawking (1971); Carr et al. (2010); Dai et al. (2009); Clesse and García-Bellido (2015); Carr et al. (2016); Clesse and García-Bellido (2017); Clesse and Garcia-Bellido (2022); Miller (2024).
@@
The numerous hypotheses regarding the existence of DM suggest that the mass of its constituents can vary across hundreds of orders of magnitude Belenchia et al. (2022). Significant time and resources, however, have been spent on searching a somewhat narrow mass range comprising so-called Weakly Interacting Massive Particles (WIMPs) Jungman et al. (1996). WIMPs have been probed through collider experiments Goodman et al. (2010) and DM/ nucleus interactions by experiments such as CDMS Akerib et al. (2003, 2006); Ahmed et al. (2009), SuperCDMS Agnese et al. (2019), XENON Angle et al. (2008, 2011); Aprile et al. (2012, 2016, 2024), and LUX Akerib et al. (2013, 2016). Other experiments have tried to detect annihilating DM indirectly through its gamma-ray emission in the Galactic Center, galaxy clusters, or other areas, e.g. MAGIC Aleksic et al. (2012); Aleksić et al. (2016), VERITAS Holder et al. (2009), HESS Aharonian et al. (2006) and Fermi Hooper and Goodenough (2011), or through cosmic rays, e.g. AMS Aguilar et al. (2007, 2013); Bergstrom et al. (2013) or Fermi-LAT Abdo et al. (2009); Ackermann et al. (2012). Despite these efforts, however, WIMPS remain undetected [^1] DAMA claims to see a periodic oscillation of a DM signal, though this result has not yet been reproduced Petriello and Zurek (2008), and so it is worthwhile to ask whether other models could better explain DM. Thus, the experimental development from the particle physics community has widened to include probing ultralight DM through (1) torsion balance experiments, such as Eöt-Wash Su et al. (1994); Schlamminger et al. (2008) and the MICROSCOPE satellite Touboul et al. (2012); Bergé et al. (2018), whose results can be interpreted as constraints on DM, (2) resonant cavity experiments, such as ADMX Du et al. (2018), (3) the Event Horizon telescope Doeleman et al. (2009); Akiyama et al. (2019), which could measure photons lensed by boson clouds and thus affect BH images Chen et al. (2023), and which could measure changes in the polarization of the electric field when the light passes through the axion background Chen et al. (2022, 2022), and (4) astrometry experiments, such as Gaia Prusti et al. (2016) and the Roman Space Telescope Akeson et al. (2019), that can meausre stochastic metric perturbations sourced by ultralight dark matter Kim (2024). It is therefore also worth asking whether high-precision GW interferometers, such as LIGO, Virgo and KAGRA Aasi et al. (2015); Acernese et al. (2015); Aso et al. (2013), can be used to search for DM, which has also been briefly reviewed in Bertone et al. (2020).
Axions and other forms of DM could be probed via their interaction with binary neutron star inspirals that could be detected in the future with LIGO, Virgo and KAGRA Hook and Huang (2018); Croon et al. (2018). The observational signature of axions on binary neutron star inspirals would be a dephasing effect induced by each object having a scalar charge, and the mass of the scalar field Hook and Huang (2018); Zhang et al. (2021). Furthermore, a variety of follow-up studies Sagunski et al. (2018); Huang et al. (2019) have been performed regarding the detectability of the effects of axions on binary neutron star inspirals, with Ref. Zhang et al. (2021) placing the first-ever constraints on the nuclear coupling of axion-like particles with masses below $10^{-11}$ eV using the detection of GWs from the binary neutron star GW170817, effectively ruling out axion decay constants of [$1.6\times 10^{16}$,$10^{18}$] GeV. Furthermore, even the very slow orbit of the Hulse-Taylor binary and other known binaries constrains the axion decay constant to be below certain values depending on the system Kumar Poddar et al. (2020) to be $f_{a}<\mathcal{O}(10^{11})$ GeV, and resonances could appear during the orbit when the orbital frequency of the binary matches that of the axion Kadota et al. (2024); Kim and Yang (2024). Finally, axions and the DM energy density can also be constrained via the evidence for stochastic GW background from pulsar timing arrays Guo et al. (2024a); Chen et al. (2024a) and also explain the mass gap in observed BHs Croon and Sakstein (2023).
@@
In particular, in this review, we would like to explore how continuous gravitational-wave (CWs) can be used as a probe of DM. CWs are quasi-monochromatic, quasi-infinite-duration signals that canonically arise from asymmetrically rotating, lumpy neutron stars. Over the last few decades, much research has focused on developing ways to probe such neutron stars, both those that are known electromagnetically and those that may only be emitting GWs (“gravitars”) Sieniawska and Bejger (2019); Tenorio et al. (2021); Riles (2023); Piccinni (2022); Miller (2023); Wette (2023). Though the signal model is simple, the unknown sky positions and the uncertain neutron-star physics (e.g. the unknown equation of state, “spin wandering” Akmal et al. (1998); Cutler (2002), changes in the magnetosphere Lyne et al. (2010), etc.) complicates these searches, and thus methods that are robust not only to non-Gaussian noise disturbances but also to theoretical uncertainties had to be developed to look for such systems Krishnan et al. (2004); Dhurandhar et al. (2008); Astone et al. (2014); Suvorova et al. (2016); Bayley et al. (2019). Until now, such searches have only yielded constraints on the maximum possible deformation of the size of known Abbott et al. (2021a, 2022a); Ashok et al. (2021, 2024); Abac et al. (2025) and unknown neutron stars Abbott et al. (2022b); Steltner et al. (2023).
It turns out that with CWs, we can not only probe neutron stars, but also a wide range of DM models. If DM particles couple to standard model ones, we could see them in GW data via their interactions with detector components, which would give rise to a narrow-band, stochastic “correlated-noise” signal Pierce et al. (2018); Guo et al. (2019); Grote and Stadnik (2019); Ismail et al. (2023). Separately, if a new boson that may or may not be DM exists and couples to gravity, and is of just the right mass, it will form clouds around rotating BHs, extracting mass and spin from them as they grow, and emitting CWs as the DM particles annihilate Brito et al. (2015). Heavy, strongly-interacting DM could induce collapses of stars that form in binaries, and as these objects orbit each other, they will emit GWs at particular frequencies distinct from if they were ordinary stars of the same mass Singh et al. (2021); Bhattacharya et al. (2024). Such ultralight particles could also interact with the material inside neutron stars, changing their moments of inertia and thus their CW emission Khmelnitsky and Rubakov (2014). CWs can be used to probe the millisecond pulsar hypothesis to the GeV excess, which, as it is more constrained, lends more strength to the WIMP DM annihilation hypothesis for the GeV excess Calore et al. (2019); Miller and Zhao (2023); Bartel and Profumo (2024).
In LABEL:fig:summary, we show a landscape of the kinds of particle DM that could be probed with current and future interferometers, both by using GWs and and through direct detection of particles interacting with the instruments. There are thus so many different kinds of DM that can be searched for by using instruments that were not strictly designed to probe the existence of DM.
@@
This review article is broken into different parts depending both on the type of DM considered and constraints on DM interactions with gravity or standard model particles. In Section II , we describe different ways that ultralight dark matter could couple to standard model particles in GW interferometers. Section III details the constraints that have been derived from searches of recent LIGO, Virgo and KAGRA data on the different ways that DM could couple to the interferometers. Section IV focuses on how macroscopic DM could transit through the interferometers, giving rise to a measurable differential acceleration of the mirrors. In Section V, we describe how scalar and vector bosons (that could be DM) form around rotating BHs, and the types of GW signals that could be emitted if these BHs are isolated or are in binary systems. Search results for annihilating scalar and vector boson cloud systems are explained in Section VI. Next, we discuss how soliton DM would impact GW signals from binary systems and neutron stars in Section VII. The collection of infamous WIMPs around BHs and in the Galactic Center are discussed in Section VIII in the context of how they can be probed with GWs. We explain how DM spikes around rotating BHs could alter GW signals in Section IX. In Section X, we describe a model of atomic DM that would give rise to “dark atoms” and “dark BHs”, and how it can be constrained with GW observations and non-observations of merging compact objects. We conclude in Section XI about the future for the burgeoning field of GW probes of particle DM.
II Ultralight dark matter interacting with the interferometers
II.1 Generic features of ultralight dark matter signal

### v4

@@ I Introduction
Our knowledge of physics has been incomplete for decades, if not for over a century Bertone and Hooper (2018). Dark energy and dark matter (DM) comprise 69% and 26% of the universe, respectively, and yet we have not been able to directly detect either. The experimental evidence for the existence of DM is overwhelming: stars orbiting the center of the spiral galaxies would have flown out of their orbits if DM did not hold them in place Freeman (1970); Rogstad and Shostak (1972); Whitehurst and Roberts (1972); Roberts and Rots (1973); anisotropies in the cosmic microwave background can result from DM potential wells in the early universe Peebles (1982); Burles and Tytler (1998); DM can seed the formation of large-scale structures visible today, which would have been impossible with ordinary matter alone Pagels and Primack (1982); Davis et al. (1982); Blumenthal et al. (1984); Davis et al. (1985); Navarro et al. (1996); and gravitational lensing of light by galaxies in between us and the light source can be explained by DM inside the galaxy Paczynski (1986); Alcock et al. (1993); Aubourg et al. (1993); Alcock et al. (2000); Lasserre (2000); Tisserand et al. (2007).
@@
While the observational evidence shows that DM exists, it does not explain what it is made of. Can it be a single new particle, such as the axion Peccei and Quinn (1977a, b); Weinberg (1978) or dark photon Galison and Manohar (1984); Holdom (1986); Pierce et al. (2018); Filippi and De Napoli (2020), or macroscopic, in the form of PBHs formed within $\sim 1$ second of the Big Bang Green and Kavanagh (2021)? Or, can particle DM exist alongside PBHs, which may comprise a portion or the totality of DM Hawking (1971); Carr et al. (2010); Dai et al. (2009); Clesse and García-Bellido (2015); Carr et al. (2016); Clesse and García-Bellido (2017); Clesse and Garcia-Bellido (2022); Miller (2024a)?
@@
The numerous hypotheses regarding the origins of DM suggest that the mass of its constituents can vary across hundreds of orders of magnitude Belenchia et al. (2022). Significant time and resources, however, have been spent on searching a somewhat narrow mass range comprising so-called Weakly Interacting Massive Particles (WIMPs) Jungman et al. (1996). WIMPs have been probed through collider experiments Goodman et al. (2010) and DM/ nucleus interactions by experiments such as CDMS Akerib et al. (2003, 2006); Ahmed et al. (2009), SuperCDMS Agnese et al. (2019), XENON Angle et al. (2008, 2011); Aprile et al. (2012, 2016, 2024), and LZ (LUX-Zeppelin) Malling et al. (2011); Akerib et al. (2013, 2016); Mount et al. (2017); Akerib et al. (2020). Other experiments have tried to detect annihilating DM indirectly through its gamma-ray emission in the Galactic Center, galaxy clusters, or other areas, e.g. MAGIC Aleksic et al. (2012); Aleksić et al. (2016), VERITAS Holder et al. (2009), HESS Aharonian et al. (2006) and Fermi Hooper and Goodenough (2011), or through cosmic rays, e.g. AMS Aguilar et al. (2007, 2013); Bergstrom et al. (2013) or Fermi-LAT Abdo et al. (2009); Ackermann et al. (2012). Despite these efforts, however, WIMPS remain undetected [^1] DAMA claims to see a periodic oscillation of a DM signal, though this result has not yet been reproduced Petriello and Zurek (2008), and so it is worthwhile to ask whether other models can better explain DM. Thus, the experimental development from the particle physics community has widened to include probing ultralight ($\ll 1$ eV) DM through (1) torsion balance experiments, such as Eöt-Wash Su et al. (1994); Schlamminger et al. (2008) and the MICROSCOPE satellite Touboul et al. (2012); Bergé et al. (2018), whose results can be interpreted as constraints on DM, (2) resonant cavity experiments, such as ADMX Du et al. (2018), (3) the Event Horizon telescope Doeleman et al. (2009); Akiyama et al. (2019), which can measure photons lensed by boson clouds and thus affect BH images Chen et al. (2023), and which can measure changes in the polarization of the electric field when the light passes through the axion background Chen et al. (2022, 2022), and (4) astrometry experiments, such as Gaia Prusti et al. (2016) and the Roman Space Telescope Akeson et al. (2019), that can measure stochastic metric perturbations sourced by ultralight dark matter Kim (2024). In light of the diversity of experiments mentioned, it is also worth asking whether high-precision GW interferometers, such as LIGO, Virgo, KAGRA and GEO600 Affeldt et al. (2014); Dooley et al. (2016); Aasi et al. (2015); Acernese et al. (2015); Aso et al. (2013), can be used to search for DM, as previously reviewed in Bertone et al. (2020); Bertone (2024).
@@
While it has been demonstrated that constraints on axions can be set using current observations of binary neutron star inspirals (GW170817) Hook and Huang (2018); Croon et al. (2018); Zhang et al. (2021) and the Hulse-Taylor binary Kumar Poddar et al. (2020), in this review, we would like to explore how continuous gravitational-wave (CWs) can be used as a probe of DM. CWs are quasi-monochromatic, quasi-infinite-duration signals that canonically arise from asymmetrically rotating, lumpy neutron stars. Over the last few decades, much research has focused on developing ways to probe such neutron stars, both those that are known electromagnetically and those that may only be emitting GWs (“gravitars”) Sieniawska and Bejger (2019); Tenorio et al. (2021); Riles (2023); Piccinni (2022); Miller (2023); Wette (2023). Though the signal model is simple, the unknown sky positions and the uncertain neutron-star physics (e.g. the unknown equation of state, “spin wandering” Akmal et al. (1998); Cutler (2002), changes in the magnetosphere Lyne et al. (2010), etc.) complicate these searches, and thus methods that are robust not only to non-Gaussian noise disturbances but also to theoretical uncertainties had to be developed to look for such systems Krishnan et al. (2004); Dhurandhar et al. (2008); Astone et al. (2014); Suvorova et al. (2016); Bayley et al. (2019). Until recently, such searches have only yielded constraints on the maximum possible deformation of the size of known Abbott et al. (2021a, 2022a); Ashok et al. (2021, 2024); Abac et al. (2025a) and unknown neutron stars Abbott et al. (2022b); Steltner et al. (2023).
In this review, we show how CW methods have been generalized to search for the following diverse set of DM models:
  1. 1.
    Ultralight bosons (scalar, vector, or tensor): Oscillating fields can couple to standard-model particles, leading to periodic modulations in the interferometer arm lengths, refractive indices, neutron-star moment of inertias, or timings of atomic clocks Khmelnitsky and Rubakov (2014); Pierce et al. (2018); Guo et al. (2019); Grote and Stadnik (2019); Ismail et al. (2023). CW methods are sensitive to these narrowband, quasi-monochromatic signals and can be applied to searches for this kind of DM in ground-based interferometers, pulsar timing arrays, or future space-based antennas. Depending on the particular interferometer, ultralight dark matter masses between $10^{-22},10^{-11}]$ eV can be probed, showing how instruments not even designed to search for DM are able to cover a wide range of masses and thus a variety of DM models.
  2. 2.
    Macroscopic dark matter objects: Compact dark objects of masses of $\mathcal{O}(1-10^{9})$ kg passing through or near GW interferometers could induce transient but quasi-periodic disturbances in the test masses Hall et al. (2018); Du et al. (2023). Searches using matched filtering can place constraints on Yukawa-type interactions or on the rate of such transient events. Even though individually, each transient would manifest as short-lived perturbation of the detector, if multiple encounters were to occur over the observation time, the resulting signature would appear as a stochastic background of transients. Potentially, CW methods could be applied to look for the periodicity or stochastic background of such transients. Moreover, networks of interferometers could help discriminate these events from terrestrial noise, providing both spatial and temporal correlations that enhance detection prospects.
  3. 3.
    Boson clouds around rotating black holes: Ultralight bosons can grow around black holes via superradiance, accelerated when the Compton wavelength of the boson matches the size of the BH, forming clouds that extract energy and momentum from the BH and emit nearly monochromatic GWs through annihilation or level transitions Arvanitaki and Dubovsky (2011); Arvanitaki et al. (2015a); Brito et al. (2015, 2017). CW searches can target these long-lived signals and test the existence of bosons in the $10^{-13},10^{-11}]$ eV mass range in ground-based GW interferometers, and between $10^{-17},10^{-15}]$ eV in future space-based GW antennas.
  4. 4.
    Soliton dark matter: Gravitational interactions can lead to the formation of solitonic cores of ultralight dark matter at the centers of galaxies Schive et al. (2014); Hui et al. (2017). Compact binaries or isolated neutron stars moving through solitons experience additional gravitational potentials that can modify the phase evolution of the GWs as binary systems inspiral, and as neutron stars spin down, respectively. For binaries, this manifests as a long-lived, cumulative dephasing effect during the inspiral Aghaie et al. (2024), while for isolated nonaxisymmetric neutron stars, frequency modulations in the emitted CWs Blas et al. (2025) are induced. In both cases, the effect is coherent over long timescales, making CW search methods particularly well suited to test these scenarios. Depending on the soliton mass and size, signatures would be detectable with pulsar timing array data (probing supermassive binary inspirals crossing galactic cores) or with next-generation terrestrial detectors such as ET and Cosmic Explorer (probing stellar-mass binaries and isolated pulsars embedded in soliton potentials).
  5. 5.
    WIMP dark matter: Even though WIMPs are not inherently wave-like, their capture in stars can induce their collapse into BHs Singh et al. (2021); Bhattacharya et al. (2024). CW searches can probe this scenario indirectly, by looking for the long-lived inspirals of BHs space-based GW interferometers. Additionally, null results from CW searches from millisecond pulsars in the Galactic Center can indirectly preference WIMP DM as responsible for the gamma-ray GeV excess emanating from the center of the Milky Way Calore et al. (2019); Miller and Zhao (2023).
  6. 6.
    Dark matter spikes around compact objects: Dense DM profiles near BHs inspiraling can generate dynamical friction as the secondary object flows through the DM spike of the first, leading to small but measurable dephasing in these long-lived inspirals Gondolo and Silk (1999); Eda et al. (2013, 2015); Kavanagh et al. (2020); Coogan et al. (2022); Cole et al. (2023a, b); Bertone (2024). While originally such effects were considered around extreme-mass ratio inspirals (EMRIs), intermediate-mass ratio inspirals (IMRIs) and ordinary binaries have been shown to allow the presence of DM spikes, and such dephasing could be measurable in ET, Cosmic Explorer or space-based GW interferometers, when inspiraling binaries will spend hours, days or years in-band. CW methods can thus test the presence and properties of such spikes.
  7. 7.
    Atomic dark matter: If DM possesses a composite structure with a dark sector analogue of electromagnetism, it could form bound states or “dark atoms” Feng et al. (2009); Kaplan et al. (2010); Buckley and DiFranzo (2018); Shandera et al. (2018). Such atomic DM can exhibit dissipative dynamics, allowing it to cool and collapse differently from standard cold DM. This can lead to the formation of compact dark objects, modify binary black hole merger populations, or even produce sub-solar mass binaries that would be invisible electromagnetically. While GW searches have constrained only the presence of such atomic DM using matched filtering searches, planetary- or asteroid-mass systems (masses $0.1M_{\odot}$) could also have been formed from atomic DM Fan et al. (2013a, b); Buckley and DiFranzo (2018), and would emit longer-duration GWs. Thus, CW searches can be applied to look for such systems Miller et al. (2022a); Abbott et al. (2022b); Miller et al. (2024).
In each of aforementioned scenarios, considering DM within the framework of CWs allows us to adapt methods originally designed to probe neutron stars to search for DM. This is especially important given the variety and complexity of the underlying physics in different DM models, as well as the non-stationary, non-Gaussian noise characteristic of GW detectors. Traditionally, matched filtering has been the primary technique for detecting GW signals from binary black hole and binary neutron star mergers, in which known templates are correlated with observed data to extract weak signals embedded in noise. However, this approach becomes less effective when dealing with exotic scenarios, such as those discussed in this review, where the signals may not conform to standard templates, or where the creation of non-traditional templates is computationally infeasible.
Importantly, we consider two general scenarios for the detection of DM using GW interferometers: (1) DM interacts either gravitationally or via couplings to standard model particles, leaving a direct imprint on interferometers, similar to traditional direct detection experiments; and (2) DM sources or modifies GW emission itself, which is then observed by interferometers.
In Fig. 1, we present a landscape of the types of particle DM that can be probed with current and future interferometers in the context of the first scenario. For each DM candidate, the figure maps out the interferometer component or astrophysical system it affects, the resulting observable, and the classes of interferometers capable of detecting the effect. Note that scenario (1) does not lead to GW signals; rather, the “signal” arises from the coupling of DM to the standard model.
In contrast, Fig. 2 summarizes the second scenario, showing how GW signals can be sourced or modified by the presence of DM. Together, these plots provide a high-level overview of the key models and observables discussed throughout this review.
[…]
Figure 2: Similar to Fig. 1. This plot shows the kinds of DM that can be probed via their generation of GWs, both now and in the future. The logic is as follows, moving from left to right: type of DM $\rightarrow$ astrophysical source affected $\rightarrow$ observable $\rightarrow$ which detector(s) the signal could be seen in. We note that the mass range of particle DM probed here ranges many orders of magnitude, from $\mathcal{O}(10^{-23}-10^{-21})$ eV with pulsar timing arrays to $\mathcal{O}(10^{-16}-10^{-13})$ eV with space-based detectors to $\mathcal{O}(10^{-13}-10^{-11})$ eV with ground-based interferometers, and includes $10^{3},10^{9}]$ GeV WIMP DM, probes of large-scale DM structure via solitons and DM spikes, and atomic DM. Matching colors imply that the source would induce those physical signatures.
@@
This review article is broken into different parts depending both on the type of DM considered and constraints on DM interactions with gravity or standard model particles. In Section II, we describe different ways that ultralight dark matter could couple to standard model particles in GW interferometers. Section III details the constraints that have been derived from searches of recent LIGO, Virgo and KAGRA data on the different ways that DM could couple to the interferometers. Section IV focuses on how macroscopic DM could transit through the interferometers, giving rise to a measurable differential acceleration of the mirrors. In Section V, we describe how scalar and vector bosons form around rotating BHs, and the types of GW signals that could be emitted if these BHs are isolated or are in binary systems. Search results for annihilating scalar and vector boson cloud systems are explained in Section VI. Next, we discuss how soliton DM would impact GW signals from binary systems and neutron stars in Section VII. The collection of WIMPs around BHs and in the Galactic Center are discussed in Section VIII in the context of how they can be probed with GWs. We explain how DM spikes around rotating BHs could alter GW signals in Section IX. In Section X, we describe a model of atomic DM that would give rise to “dark atoms” and “dark BHs”, and how it can be constrained with GW observations and non-observations of merging compact objects. We conclude in Section XI about the future for the burgeoning field of GW probes of particle DM.
Note that we use natural units throughout this review: $c=\hbar=\epsilon_{0}=1$.
II Ultralight DM interacting with GW interferometers
Numerous models for DM exist throughout the literature, and one review cannot possibly cover all of them. Instead, in this section, we focus specifically on ultralight dark matter models that have been shown to cause observable signals in GW interferometers. Over the last decade, the impact of more and more DM models on ground- and space-based instruments has been considering, allowing the field of “direct DM searches with GW interferometers” to grow. The models we present here should be taken as the minimum number of ways in which DM could affect GW interferometers. In Section II.1, we outline some general properties of all ultralight dark matter considered in this review, and then delve into four types of DM that couple to the standard model: axions (Section II.2), dilatons (Section II.3), vector, dark photon DM (Section II.4), tensor DM (Section II.5). We end with a scalar DM field that couples only to gravity (Section II.6).
II.1 Generic features of ultralight dark matter signals